|
Документ взят из кэша поисковой машины. Адрес
оригинального документа
: http://www.stsci.edu/smov/notes/0304smov.html
Дата изменения: Tue Mar 11 23:59:22 1997 Дата индексирования: Thu Feb 27 23:54:35 2014 Кодировка: Поисковые слова: http www.badastronomy.com phpbb index.php |
STATUS REPORTS
1. Observatory Status
MOSES status reports:
- All observatory subsystems all nominal
2. WFPC2 Status
- Engineering status is nominal. The scheduled DECON
was performed last night and the instrument is re-cooled
to -88C.
- Science status - nominal. The latest UV throughput
measurements are reducing the error bars enough to
conclude that there is small amount of excess
contamination buildup. The UV loss rate is now reported at
1.1%/day, +/- 0.3%), compared to pre-SM2 rate of 0.7%/day.
This amount is not considered serious enough to resume
bright-Earth avoidance (BEA) or to prevent the resumption of
Cycle 6 science and ERO on March 6. Howver WFPC2 has
requested an earlier decontamination, in two weeks instead
of four. (The calendar changes have already been requested
and made.)
3. NICMOS Status
- Engineering status - Yesterday's discussions produced an
approach to today's Pre-Alignment Check-Out (7134, which
called for the Pre-postioning of the PAM during the day
Monday, after which NICMOS would be brought down to BOOT
mode where it is not susceptible to SAA-induced SUSPENDs.
These activities were performed per the plan. Then this
morning, TDRSS links were acquired/extended which allowed
reconfiguring the instrument to Operate in preparation
for the Pre-Alignmnent Check-Out via stored commanding.
This is thought to be a relatively robust plan that
reduces the chances of losing the activity to an
SAA-induced anomaly. However, the consensus was that,
should another anomaly occur while in OPERATE prior to the
start of the activity later in the morning, NICMOS would be
left in SUSPEND mode rather than recovered under time
pressure.
- Science status - Further discussions yesterday also resulted
in changes to the Coarse Corrector Alignment Proposal
(7041), scheduled in SMS069, which succeeded in moving
several visits into SAA-free zones. (The proposal had been
scheduled in this calendar before the SAA susceptibilities
were known.) This was achieved at the expense of some
previously scheduled WFPC2 Cycle 6 science.
- It was also decided to remove the SAA Contour Test (7036)
from the SMSs 072 and 079, and to redesign it for much
shallower SAA penetration as a means of mapping
NICMOS-specific SAA contours.
- Other downstream NICMOS proposals are being reviewed for
any other changes that may be made necessary by the SAA
susceptibilities.
4. STIS Status
- Engineering Status - Nominal - The MAMA LVPSs remain
off indefinitely, pending further investigation by BASD.
The BASD life-testing of the Band 2 spare, mentioned in
yesterday's notes, is deemed by the PI to be a lower
priority than the on-going LV/HV health and safety analysis
that are needed to proceed operationally with MAMA SMOV.
- Science Status - As reported yesterday, the first STIS
external observation, a star field viewed by the CCD
for the purpose of STIS/FGS alignment (7065) looks quite
good. In addition, the data from the first STIS CCD UV
monitor (7063) has been received and looks nominal.
5. FGS status - no activity
- The first FGS1R activity (7026, visit 1) was successfully
executed yesterday morning. HDOS and STScI reported that
the "miss distance" was approximately 7 arcseconds, which
is "not bad." The PMT counts are as predicted. The
y-axis shows a 1-mm decenter and the x-axis a 0.1-mm
decenter. A commanding problem prevented the ITS
position test; this has already been corrected for
performance in a later visit of this proposal.
Analysis continues prior to next week's scheduled
FF3 moves and S-curve measurements.
6. FOC/COSTAR status
The FOC non-COSTAR baseline throughput measurements (7012)
were successfully executed. The data have been received and
archived. The images and PSF look normal (pre-COSTAR), though the
data have not yet been analyzed in detail.
PLANNING REPORTS
1. Status of Upcoming SMSs
- SMS062 - Currently executing.
- SMS065 - Signed off, ready for uplink.
- SMS069 - Final Mode products are out for review. This
SMS was rebuilt last night to include the NICMOS changes
mentioned above.
- SMS072 - Rebuilt without the NICMOS 7036 mentioned above.
- SMS076 - Built and initials out for review
- SMS079 - Initial build complete. This SMS currently
contains the 1st part of the MAMA HV turn-on,
per the nominal plan.
2. Replans - NICMOS 7041 was modified withing SMS069. NICMOS 7036
was removed, for rework, from the SMS072.
OPERATIONS STATUS
1. OPUS - Nothing new to report.
2. Archive - A problem in getting STIS reference files out of the
archive is being worked.